The outstanding capital stock of Edna Millay Corporation consists of 2,000 shares of $100 par value, 8% preferred, and 5,000 shares of $50 par value common.

Assuming that the company has retained earnings of $90,000, all of which is to be paid out in dividends, and that preferred dividends were not paid during the 2 years preceding the current year, state how much each class of stock should receive under each of the following conditions.

(a) The preferred stock is noncumulative and nonparticipating. (Round answers to 0 decimal places, e.g. $38,487.)

Preferred

Common

$The outstanding capital stock of Edna Millay Corpo $The outstanding capital stock of Edna Millay Corpo


(b) The preferred stock is cumulative and nonparticipating. (Round answers to 0 decimal places, e.g. $38,487.)

Preferred

Common

$The outstanding capital stock of Edna Millay Corpo $The outstanding capital stock of Edna Millay Corpo


(c) The preferred stock is cumulative and participating. (Round the rate of participation to 4 decimal places, e.g.1.4278%. Round answers to 0 decimal places, e.g. $38,487.)

Preferred

Common

$The outstanding capital stock of Edna Millay Corpo $The outstanding capital stock of Edna Millay Corpo
Answer #1
Concepts and reason

Stock: An equal part or a portion of the company’s capital which indicates ownership of the holder in the company as well as states that the shareholder has a claim in the profits and the assets of the company is referred to as a share. A bundle of fully paid shares is referred to as stock. The two main categories of stock are preferred stock and common stock.

Fundamentals

Common stock: Common stock is the stock issued by a company to the people having voting rights. It is termed as stockholders’ equity account. It is a form of corporate equity ownership. It is one of the forms of securities that are issued to the common shareholders. They are entitled to dividend and repayment of capital after the payment is made to preference shareholders. It will be reported in the balance sheet as liabilities.

Preferred stock: Preferred stock is the stock issued by a company, indicating that the holders of the preferred stock are given priority while issuing dividend. They are completely settled and in case of arrears, the dividend or any amount is settled in the future year.

Cumulative preferred stock: Cumulative preferred stock indicates that when the payment of dividend has not been made to appropriate shareholders, it should be done in the near future, and the shareholders must be considered as the first priority during the next payment of dividend or if any inter-dividend or profit is earned.

Dividends: Dividends are the divisible profits of a company that are issued to the shareholder for the portion of share they have purchased. It should be through the way of purchase in the company. It is an expense for the company.

1)

Calculate the total dividends paid:

Total
$16,000
Particulars
Current Dividends Paid to Preferred
Add: Paid to Common Shareholders
($90,000-$16,000)
Total Divide

Therefore, the total dividend paid is $90,000.

Working notes:

Determine the amount of preferred current dividend paid:

PreferredCurrentDividend=NumberofShares×FaceValue×RateofDividend=2,000shares×$100×8%=$200,000×8%=$16,000\begin{array}{c}\\{\rm{Preferred Current Dividend = Number of Shares }} \times {\rm{ Face Value }} \times {\rm{ Rate of Dividend}}\\\\{\rm{ = 2,000shares }} \times {\rm{ \$ 100 }} \times {\rm{ 8\% }}\\\\{\rm{ = \$ 200,000}} \times {\rm{ 8\% }}\\\\{\rm{ = \$ 16,000}}\\\end{array}

Therefore, the amount of preferred current dividend paid is $16,000.

2)

Calculate the total dividends paid for non-participating and non-cumulative preferred shareholders:

Total
$32,000
Particulars
Preferred Common
Preferred dividend in arrears
$32,000
|(16,000*2)
Current Dividends Paid to Prefer

Therefore, the total dividend paid to preferred shareholders is $48,000 and a common shareholder is $42,000.

3)

Calculate the total dividends paid for participating and cumulative preferred stock:

Preferred Common
Total
$32.000
$32,000
Particulars
Preferred dividend in arrears
(16,000*2)
Current Dividends Paid to
Preferr

Therefore, the total dividends paid to common shareholders are $32,222 and to a preferred shareholder is $57,778.

Working notes:

Calculation is given below:

12
Preferred
Common
Total
=16000*2
32000
Particulars
Preferred dividend in arrears
(16,000*2)
Current Dividends Paid to
Prefe

Determine the participating dividend rate:

PartcipatingDividendRate=RemainingbalanceofdividendTotalamount=$22,000$450,000=4.888%\begin{array}{c}\\{\rm{Partcipating Dividend Rate = }}\frac{{{\rm{Remaining balance of dividend }}}}{{{\rm{Total amount}}}}\\\\{\rm{ = }}\frac{{{\rm{\$ 22,000}}}}{{{\rm{\$ 450,000}}}}\\\\{\rm{ = 4}}{\rm{.888\% }}\\\end{array}

Therefore, the participating dividend rate is 4.888%.

TotalAmount=[(Numberpreferredshares×Facevalue)+(Numbercommonshares×Facevalue)]=(2,000shares×$100)+(5,000shares×$50)=$200,000+$250,000=$450,000\begin{array}{c}\\{\rm{Total Amount = }}\left[ \begin{array}{l}\\\left( {{\rm{Number preferred shares \times Face value}}} \right)\\\\{\rm{ + }}\left( {{\rm{Number common shares \times Face value}}} \right)\\\end{array} \right]\\\\{\rm{ = }}\left( {{\rm{2,000shares \times \$ 100}}} \right){\rm{ + }}\left( {{\rm{5,000 shares \times \$ 50}}} \right)\\\\{\rm{ = \$ 200,000 + \$ 250,000}}\\\\{\rm{ = \$ 450,000}}\\\end{array}

Therefore, the total amount is $450,000.

Ans: Part 1

Particulars
Current Dividends Paid to Preferred
Add: Paid to Common Shareholders
Total Dividends Paid
Preferred Common
$16,00 Particulars
Current Dividends Paid to Preferred
Add: Paid to Common Shareholders
Total Dividends Paid
Preferred Common
$15,00

Part 2

Preferred Common
$32,000
Total
$32,000
Particulars
Preferred dividend in arrears
Current Dividends Paid to
Preferred
Add: Pai

Part 3

Preferred
Common
Total
$32,000
$32,000
Particulars
Preferred dividend in
arrears
Current Dividends Paid to
Preferred
Add: Pai